# Frobenius pull-back of parabolic bundles and dormant opers

### v1

@@ 3.1. Log flat bundles and higher-level $\mathcal{D}$-modules
Let us fix a positive integer $N$. For simplicity, we use the notation “$\mathcal{D}^{(N-1)}$” to denote the sheaf $\mathcal{D}^{(N-1)}_{X^{\mathrm{log}}/S}$. For each integer $j$, we shall write $\mathcal{D}^{(N-1)}_{<j}$ for the subsheaf of $\mathcal{D}^{(N-1)}$ consisting of logarithmic differential operators of order $<j$.
@@ 3.2. Monodromy operators
Next, let $(\mathcal{F},\nabla)$ be a $\mathcal{D}^{(N-1)}$-module, and choose $i\in\{1,\cdots,r\}$. Note that there exists a canonical isomorphism of $\mathcal{O}_{S}$-algebras $\sigma_{i}^{*}({{}^{L}}\mathcal{D}^{(N-1)})\xrightarrow{\sim}\mathcal{B}_{S}$ such that the local basis $\{\partial^{\langle j\rangle}\}_{j\in\mathbb{Z}_{\geq 0}}$ of ${{}^{L}}\mathcal{D}^{(N-1)}$ associated to any logarithmic coordinate around a point in $\mathrm{Im}(\sigma_{i})$ is mapped to $\{\partial_{\mathcal{B}}^{\langle j\rangle}\}_{j\in\mathbb{Z}_{\geq 0}}$ (cf.  [Wak4, Lemma 2.2.2 and § 6.1]). The $\mathcal{D}^{(N-1)}$-module structure $\nabla$ induces an $\sigma_{i}^{*}({{}^{L}}\mathcal{D}^{(N-1)})$-action $\sigma_{i}^{*}\nabla$ on $\sigma_{i}^{*}\mathcal{F}$; it gives the composite
@@ 3.3. Parabolic $p^{N}$-flat bundles
Let $\mathscr{F}:=(\mathcal{F},\vec{\mathpzc{f}})$ be a quasi-parabolic bundle on $\mathscr{X}$. A $\mathcal{D}^{(N-1)}$-module structure on $\mathscr{F}$ is a $\mathcal{D}^{(N-1)}$-module structure $\nabla$ on $\mathcal{F}$ such that $\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})$ is closed under the $\mathcal{D}^{(N-1)}$-action $\nabla$ for every $i=1,\cdots,r$ and $j=0,\cdots,m_{i}$.
Let $\mathscr{F}:=(\mathcal{F},\vec{\mathpzc{f}})$ be a quasi-parabolic bundle on $\mathscr{X}$ and $(m_{1},\cdots,m_{r})$ the length of $\vec{\mathpzc{f}}=(\mathpzc{f}_{1},\cdots,\mathpzc{f}_{r})$. Also, let $\nabla$ be a $\mathcal{D}^{(N-1)}$-module structure on $\mathscr{F}$. Since $\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})$ (for each $i=1,\cdots,r$ and $j=1,\cdots m_{i}$) is closed under $\nabla$, we see that $\nabla$ induces a $\mathcal{B}_{S}$-module structure
$\mu_{i}^{[j]}(\nabla):\mathcal{B}_{S}\rightarrow\mathcal{E}nd_{\mathcal{O}_{S}}(\mathrm{Ker}(\mathpzc{f}_{i}^{[j]}))$ (3.9)
on $\mathrm{Ker}(\mathpzc{f}_{i}^{[j]})\left(\cong\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j-1]})/\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})\right)$; this corresponds to
$\mu_{i}^{[j]}(\nabla)^{\langle\bullet\rangle}:=(\mu_{i}^{[j]}(\nabla)^{\langle 1\rangle},\mu_{i}^{[j]}(\nabla)^{\langle p\rangle},\cdots,\mu_{i}^{[j]}(\nabla)^{\langle p^{N-1}\rangle})\in\mathrm{End}_{\mathcal{O}_{S}}(\mathrm{Ker}(\mathpzc{f}_{i}^{[j]}))^{\oplus N}.$ (3.10)
@@
    $\mathscr{F}_{\flat}:=(\mathcal{F},\nabla,\vec{\mathpzc{f}},\vec{a})$ (3.12)
@@
    • $(\mathcal{F},\vec{\mathpzc{f}},\vec{a})$ denotes a parabolic bundle on $\mathscr{X}$ such that $\vec{a}$ belongs to $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$ for some $(m_{1},\cdots,m_{r})\in\mathbb{Z}_{>0}^{\times r}$ (in particular, $(m_{1},\cdots,m_{r})$ coincides with the length of $\vec{\mathpzc{f}}$);
    • $\nabla$ denotes a $\mathcal{D}^{(N-1)}$-module structure on $(\mathcal{F},\vec{\mathpzc{f}})$ with $\psi_{(\mathcal{F},\nabla)}=0$ (resp., $(\psi_{(\mathcal{F},\nabla)})^{p-1}=0$).
    Then, we say that $\mathscr{F}_{\flat}$ is a parabolic $p^{N}$-flat (resp., parabolic $p^{N}$-nilpotent) bundle on $\mathscr{X}$ if, for any $i=1,\cdots,r$, $j=1,\cdots,m_{i}$, and $s=0,\cdots,N-1$, the element $\mu_{i}^{[j]}(\nabla)^{\langle p^{s}\rangle}\in\mathrm{End}_{\mathcal{O}_{S}}(\mathrm{Ker}(\mathpzc{f}_{i}^{[j]}))$ coincides with the multiplication by $(-a_{i}^{[j]})_{[s]}\in\mathbb{F}_{p}$.
@@
    Let $\mathscr{F}_{\flat}:=(\mathcal{F},\nabla,\vec{\mathpzc{f}},\vec{a})$ be a parabolic $p^{N}$-flat bundle on $\mathscr{X}$. By a parabolic $p^{N}$-flat subbundle of $\mathscr{F}_{\flat}$, we mean a parabolic $p^{N}$-flat bundle of the form $\mathscr{E}_{\flat}:=(\mathcal{E},\nabla|_{\mathcal{E}},\vec{\mathpzc{f}}_{\mathcal{E}},\vec{a})$, where $(\mathcal{E},\vec{\mathpzc{f}}_{\mathcal{E}},\vec{a})$ is a parabolic subbundle of $\mathscr{F}:=(\mathcal{F},\vec{\mathpzc{f}},\vec{a})$ such that $\mathcal{E}$ is closed under the $\mathcal{D}^{(N-1)}$-action $\nabla$, and $\nabla|_{\mathcal{E}}$ denotes the resulting $\mathcal{D}^{(N-1)}$-action induced by $\nabla$.
@@ 3.4. Exponent of a $p^{N}$-flat bundle
over $S$, where $t$ denotes a formal parameter; it has a closed immersion $\sigma_{\circ}:S\hookrightarrow U$ determined by “$t=0$”. The pair
@@
for some $d^{[1]},\cdots,d^{[n]}\in\mathbb{Z}/p^{N}\mathbb{Z}$; the multiset $[d^{[1]},\cdots,d^{[n]}]$ depends only on the isomorphism class of $(\mathcal{F},\nabla)$. Even in the case where $S$ is an arbitrary connected $k$-scheme (that is not necessarily the spectrum of a local ring), we can find a well-defined multiset of integers
@@ 4.1. Parabolic Frobenius pull-back in a local situation I
where ¯$\overline{\nabla}_{0}$ denotes the $\mathcal{D}^{(N-1)}$-module structure on $\mathcal{O}_{U}/\mathcal{O}_{U}((a_{\circ}^{[j]}-p^{N})\sigma_{\circ})$ induced by $\nabla_{0}$, the first arrow is obtained by extending $F^{(N)*}_{U/S}\varpi_{\mathscr{E}}^{[j]}$, and the second arrow arises from the natural quotient $(\mathcal{O}_{U},\nabla_{0})\twoheadrightarrow(\mathcal{O}_{U}/\mathcal{O}_{U}((a_{\circ}^{[j]}-p^{N})\sigma_{\circ}),\overline{\nabla}_{0})$. Hence, the kernel of this composite determines a $\mathcal{D}^{(N-1)}$-module $(\mathrm{Ker}(\widetilde{\varpi}_{\mathscr{E}}^{[j]}),\nabla^{[j]}_{\mathrm{Ker}})$. By applying this construction to various $j$’s and taking their intersection, we obtain a $\mathcal{D}^{(N-1)}$-module
$\mathscr{E}^{F}=(\mathcal{E}^{F},\nabla^{F}):=\bigcap_{j=1}^{m_{\circ}}(\mathrm{Ker}(\widetilde{\varpi}_{\mathscr{E}}^{[j]}),\nabla^{[j]}_{\mathrm{Ker}}).$ (4.4)
@@
In particular, we have $\sigma_{\circ}^{*}\mathcal{H}=\bigoplus_{j^{\prime}=1}^{m_{\circ}}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}$. For each $j=1,\cdots,m_{\circ}$, denote by $\mathpzc{h}_{\circ}^{[j]}$ the projection $\mathcal{H}_{\circ}^{[j]}\twoheadrightarrow\mathcal{H}_{\circ}^{[j-1]}$ given by forgetting the last direct summand. The triple
$\mathscr{H}:=(\mathcal{H},\mathpzc{h}_{\circ},a_{\circ}/p^{N}),$ (4.10)
where $\mathpzc{h}_{\circ}:=\{\mathpzc{h}_{\circ}^{[j]}\}_{j=1}^{m_{\circ}}$, specifies a parabolic bundle on $\mathscr{U}^{(N)}$. Since the assertion is of local nature, we can suppose that there exists an isomorphism $\tau:\mathscr{E}\xrightarrow{\sim}\mathscr{H}$. It follows that the problem can be reduced to the case of $\mathscr{E}=\mathscr{H}$.
@@
[…]
@@ 4.2. Parabolic Frobenius pull-back in a local situation II
Next, we give an inverse direction of the above construction. Let $\mathscr{F}_{\flat}:=(\mathcal{F},\nabla,\mathpzc{f}_{\circ},a_{0})$ (where $\mathscr{F}:=(\mathcal{F},\mathpzc{f}_{\circ})$, $\mathpzc{f}_{\circ}:=\{\mathpzc{f}_{\circ}^{[j]}\}_{j=1}^{m_{\circ}}$) be a parabolic $p^{N}$-flat bundle on $\mathscr{U}$ classified by $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}$. We shall set $\mathcal{F}^{\nabla}:=\mathcal{S}ol(\nabla)$, which specifies a rank-$n$ vector bundle on $U^{(N)}$. For each $j=0,\cdots,m_{\circ}$, $\nabla$ restricts to a $\mathcal{D}^{(N-1)}$-module structure $\nabla^{[j]}$ on $\mathrm{Ker}(\varpi_{\mathscr{F}}^{[j]})(-(a_{\circ}^{[j]}+1)D_{\circ})$. Then, $\mathcal{S}ol(\nabla^{[j]})$ specifies a vector bundle on $U^{(N)}$ equipped with a natural inclusion $\mathcal{S}ol(\nabla^{[j]})\hookrightarrow\mathcal{S}ol(\nabla)$, which becomes an isomorphism when restricted to $U^{(N)}\setminus\mathrm{Im}(\sigma_{\circ}^{(N)})$. The $\mathcal{O}_{S}$-modules $\mathcal{F}^{\nabla[j]}_{\circ}:=\sigma_{\circ}^{(N)*}(\mathcal{S}ol(\nabla)/\mathcal{S}ol(\nabla^{[j]}))$ fit into a sequence of surjections
$\mathpzc{f}_{\circ}^{\nabla}:\sigma_{\circ}^{(N)*}\mathcal{F}^{\nabla}=\mathcal{F}_{\circ}^{\nabla[m_{\circ}]}\xrightarrow{\mathpzc{f}_{\circ}^{\nabla[m_{\circ}]}}\mathcal{F}_{\circ}^{\nabla[m_{\circ}-1]}\xrightarrow{\mathpzc{f}_{\circ}^{\nabla[m_{\circ}-1]}}\cdots\xrightarrow{\mathpzc{f}_{\circ}^{\nabla[2]}}\mathcal{F}_{\circ}^{\nabla[1]}\xrightarrow{\mathpzc{f}_{\circ}^{\nabla[1]}}\mathcal{F}_{\circ}^{\nabla[0]}=0.$ (4.13)
@@
    $\mathscr{F}^{\nabla}:=(\mathcal{F}^{\nabla},\mathpzc{f}^{\nabla}_{\circ},a_{\circ}/p^{N})$ (4.14)
@@
    that are functorial with respect to $\mathscr{E}$ and $\mathscr{F}_{\flat}$, respectively. Moreover, when restricted to $U\setminus\mathrm{Im}(\sigma_{\circ})$, these isomorphisms coincide with the natural isomorphisms $\mathcal{E}\xrightarrow{\sim}(F^{(N)*}_{U/S}\mathcal{E})^{\nabla}$ and $F^{(N)*}_{U/S}(\mathcal{F}^{\nabla})\xrightarrow{\sim}\mathcal{F}$, respectively.
@@
First, we shall prove assertion (i). Denote by $\ell_{\circ}:=(\ell_{\circ}^{[1]},\cdots,\ell_{\circ}^{[m_{\circ}]})$ the type of $\mathpzc{f}_{\circ}$. The assertion is of local nature, so $\mathscr{F}:=(\mathcal{F},\nabla)$ may be identified with a direct sum of $\mathscr{O}_{d,\flat}$’s (cf. the discussion around (3.19)). Since $(\mathcal{F},\nabla)$ has parabolic weights $a_{0}$, the equality (3.18) implies that the problem is reduced to the situation where $\mathscr{F}$ satisfies the following conditions:
@@
  • For each $j=0,\cdots,m_{\circ}$, the equality $\mathcal{F}_{\circ}^{[j]}=\bigoplus_{j^{\prime}=1}^{j}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}$ holds. Also, the projection $\mathpzc{f}_{\circ}^{[j]}$ coincides with the projection $\mathcal{F}_{\circ}^{[j]}$ given by forgetting the last direct summand.
@@
($j=0,\cdots,m_{\circ}$), where $D^{(N)}_{\circ}:=\mathrm{Im}(\sigma_{\circ}^{(N)})$. Hence, the equality $\mathcal{F}^{\nabla[j]}_{\circ}=\bigoplus_{j^{\prime}=1}^{j}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}$ holds and $\mathpzc{f}_{\circ}^{\nabla[j]}:\mathcal{F}^{\nabla[j]}_{\circ}\twoheadrightarrow\mathcal{F}^{\nabla[j-1]}_{\circ}$ coincides the projection given by forgetting the last direct summand. Thus, $\mathscr{F}^{\nabla}$ specifies an object in $\mathcal{B}un_{\mathscr{U}^{(N)},a_{0}/p^{N}}$, and this completes the proof of assertion (i).
@@
Since the formation of $\mathscr{F}_{\flat}\mapsto\mathscr{F}^{\nabla}$ is verified to be functorial, this operation yields a functor $\mathcal{D}^{(N)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}\rightarrow\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}$.

### v2

@@ 3.1. Log flat bundles and higher-level $\mathcal{D}$-modules
Let us fix a positive integer $N$. For simplicity, we use the notation $\mathcal{D}^{(N-1)}$” to denote the sheaf $\mathcal{D}^{(N-1)}_{X^{\mathrm{log}}/S}$. For each integer $j$, we shall write $\mathcal{D}^{(N-1)}_{<j}$ for the subsheaf of $\mathcal{D}^{(N-1)}$ consisting of logarithmic differential operators of order <j<j.
@@ 3.2. Monodromy operators
[…]
@@ 3.3. Parabolic $p^{N}$-flat bundles
Let $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}})$ be a quasi-parabolic bundle on $\mathscr{X}$. A $\mathcal{D}^{(N-1)}$-module structure on $\mathscr{F}$ is a $\mathcal{D}^{(N-1)}$-module structure $\nabla$ on $\mathcal{F}$ such that $\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})$ is closed under the $\mathcal{D}^{(N-1)}$-action $\nabla$ for every $i=1,\cdots,r$ and $j=0,\cdots,m_{i}$.
Let $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}})$ be a quasi-parabolic bundle on $\mathscr{X}$ and $(m_{1},\cdots,m_{r})$ the length of $\vec{\mathfrak{f}}=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$. Also, let $\nabla$ be a $\mathcal{D}^{(N-1)}$-module structure on $\mathscr{F}$. Since $\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})$ (for each $i=1,\cdots,r$ and $j=1,\cdots m_{i}$) is closed under $\nabla$, we see that $\nabla$ induces a $\mathcal{B}_{S}$-module structure
$\mu_{i}^{[j]}(\nabla):\mathcal{B}_{S}\rightarrow\mathcal{E}nd_{\mathcal{O}_{S}}(\mathrm{Ker}(\mathfrak{f}_{i}^{[j]}))$ (3.9)
on $\mathrm{Ker}(\mathfrak{f}_{i}^{[j]})\left(\cong\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j-1]})/\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})\right)$; this corresponds to
$\mu_{i}^{[j]}(\nabla)^{\langle\bullet\rangle}:=(\mu_{i}^{[j]}(\nabla)^{\langle 1\rangle},\mu_{i}^{[j]}(\nabla)^{\langle p\rangle},\cdots,\mu_{i}^{[j]}(\nabla)^{\langle p^{N-1}\rangle})\in\mathrm{End}_{\mathcal{O}_{S}}(\mathrm{Ker}(\mathfrak{f}_{i}^{[j]}))^{\oplus N}.$ (3.10)
@@
    $\mathscr{F}_{\flat}:=(\mathcal{F},\nabla,\vec{\mathfrak{f}},\vec{a})$ (3.12)
@@
    • $(\mathcal{F},\vec{\mathfrak{f}},\vec{a})$ denotes a parabolic bundle on $\mathscr{X}$ such that $\vec{a}$ belongs to $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$ for some $(m_{1},\cdots,m_{r})\in\mathbb{Z}_{>0}^{\times r}$ (in particular, $(m_{1},\cdots,m_{r})$ coincides with the length of $\vec{\mathfrak{f}}$);
    • $\nabla$ denotes a $\mathcal{D}^{(N-1)}$-module structure on $(\mathcal{F},\vec{\mathfrak{f}})$ with $\psi_{(\mathcal{F},\nabla)}=0$ (resp., $(\psi_{(\mathcal{F},\nabla)})^{p-1}=0$).
    Then, we say that $\mathscr{F}_{\flat}$ is a parabolic $p^{N}$-flat (resp., parabolic $p^{N}$-nilpotent) bundle on $\mathscr{X}$ if, for any $i=1,\cdots,r$, $j=1,\cdots,m_{i}$, and $s=0,\cdots,N-1$, the element $\mu_{i}^{[j]}(\nabla)^{\langle p^{s}\rangle}\in\mathrm{End}_{\mathcal{O}_{S}}(\mathrm{Ker}(\mathfrak{f}_{i}^{[j]}))$ coincides with the multiplication by $(-a_{i}^{[j]})_{[s]}\in\mathbb{F}_{p}$.
@@
    Let $\mathscr{F}_{\flat}:=(\mathcal{F},\nabla,\vec{\mathfrak{f}},\vec{a})$ be a parabolic $p^{N}$-flat bundle on $\mathscr{X}$. By a parabolic $p^{N}$-flat subbundle of $\mathscr{F}_{\flat}$, we mean a parabolic $p^{N}$-flat bundle of the form $\mathscr{E}_{\flat}:=(\mathcal{E},\nabla|_{\mathcal{E}},\vec{\mathfrak{f}}_{\mathcal{E}},\vec{a})$, where $(\mathcal{E},\vec{\mathfrak{f}}_{\mathcal{E}},\vec{a})$ is a parabolic subbundle of $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{a})$ such that $\mathcal{E}$ is closed under the $\mathcal{D}^{(N-1)}$-action $\nabla$, and $\nabla|_{\mathcal{E}}$ denotes the resulting $\mathcal{D}^{(N-1)}$-action induced by $\nabla$.
@@ 3.4. Exponent of a $p^{N}$-flat bundle
over $S$, where $t$ denotes a formal parameter; it has a closed immersion $\sigma_{\circ}:S\hookrightarrow U$ determined by “t=0t=0”. The pair
@@
for some $d^{[1]},\cdots,d^{[n]}\in\mathbb{Z}/p^{N}\mathbb{Z}$; the multiset $d^{[1]},\cdots,d^{[n]}]$ depends only on the isomorphism class of $(\mathcal{F},\nabla)$. Even in the case where $S$ is an arbitrary connected $k$-scheme (that is not necessarily the spectrum of a local ring), we can find a well-defined multiset of integers
@@ 4.1. Parabolic Frobenius pull-back in a local situation I
[…]
@@
In particular, we have $\sigma_{\circ}^{*}\mathcal{H}=\bigoplus_{j^{\prime}=1}^{m_{\circ}}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}$. For each $j=1,\cdots,m_{\circ}$, denote by $\mathfrak{h}_{\circ}^{[j]}$ the projection $\mathcal{H}_{\circ}^{[j]}\twoheadrightarrow\mathcal{H}_{\circ}^{[j-1]}$ given by forgetting the last direct summand. The triple
$\mathscr{H}:=(\mathcal{H},\mathfrak{h}_{\circ},a_{\circ}/p^{N}),$ (4.10)
where $\mathfrak{h}_{\circ}:=\{\mathfrak{h}_{\circ}^{[j]}\}_{j=1}^{m_{\circ}}$, specifies a parabolic bundle on $\mathscr{U}^{(N)}$. Since the assertion is of local nature, we can suppose that there exists an isomorphism $\tau:\mathscr{E}\xrightarrow{\sim}\mathscr{H}$. It follows that the problem can be reduced to the case of $\mathscr{E}=\mathscr{H}$.
@@
($j=0,\cdots,m_{\circ}$), and the projection $\mathfrak{h}_{\circ}^{F[j]}$ is given by forgetting the last direct summand. Thus, $(\mathcal{E}^{F},\mathfrak{h}_{\circ}^{F})$ forms a quasi-parabolic bundle on $\mathscr{U}$. Moreover, since $(\mathcal{O}_{U}(a_{\circ}^{[j]}D_{\oslash}),\nabla_{0})\cong(\mathcal{O}_{U},\nabla_{a_{\circ}^{[j]}})$, the description of $\nabla_{\mathrm{Ker}}^{[j]}$ obtained in (4.11) and the equality (3.18) together imply that $\mathscr{H}^{F}_{\flat}:=(\mathcal{E}^{F},\nabla^{F},\mathfrak{h}_{\circ}^{F},a_{\circ})$ belongs to $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}$. This completes the proof of the first assertion.
@@ 4.2. Parabolic Frobenius pull-back in a local situation II
Next, we give an inverse direction of the above construction. Let $\mathscr{F}_{\flat}:=(\mathcal{F},\nabla,\mathfrak{f}_{\circ},a_{0})$ (where $\mathscr{F}:=(\mathcal{F},\mathfrak{f}_{\circ})$, $\mathfrak{f}_{\circ}:=\{\mathfrak{f}_{\circ}^{[j]}\}_{j=1}^{m_{\circ}}$) be a parabolic $p^{N}$-flat bundle on $\mathscr{U}$ classified by $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}$. We shall set $\mathcal{F}^{\nabla}:=\mathcal{S}ol(\nabla)$, which specifies a rank-$n$ vector bundle on $U^{(N)}$. For each $j=0,\cdots,m_{\circ}$, $\nabla$ restricts to a $\mathcal{D}^{(N-1)}$-module structure $\nabla^{[j]}$ on $\mathrm{Ker}(\varpi_{\mathscr{F}}^{[j]})(-(a_{\circ}^{[j]}+1)D_{\circ})$. Then, $\mathcal{S}ol(\nabla^{[j]})$ specifies a vector bundle on $U^{(N)}$ equipped with a natural inclusion $\mathcal{S}ol(\nabla^{[j]})\hookrightarrow\mathcal{S}ol(\nabla)$, which becomes an isomorphism when restricted to $U^{(N)}\setminus\mathrm{Im}(\sigma_{\circ}^{(N)})$. The $\mathcal{O}_{S}$-modules $\mathcal{F}^{\nabla[j]}_{\circ}:=\sigma_{\circ}^{(N)*}(\mathcal{S}ol(\nabla)/\mathcal{S}ol(\nabla^{[j]}))$ fit into a sequence of surjections
$\mathfrak{f}_{\circ}^{\nabla}:\sigma_{\circ}^{(N)*}\mathcal{F}^{\nabla}=\mathcal{F}_{\circ}^{\nabla[m_{\circ}]}\xrightarrow{\mathfrak{f}_{\circ}^{\nabla[m_{\circ}]}}\mathcal{F}_{\circ}^{\nabla[m_{\circ}-1]}\xrightarrow{\mathfrak{f}_{\circ}^{\nabla[m_{\circ}-1]}}\cdots\xrightarrow{\mathfrak{f}_{\circ}^{\nabla[2]}}\mathcal{F}_{\circ}^{\nabla[1]}\xrightarrow{\mathfrak{f}_{\circ}^{\nabla[1]}}\mathcal{F}_{\circ}^{\nabla[0]}=0.$ (4.13)
@@
    $\mathscr{F}^{\nabla}:=(\mathcal{F}^{\nabla},\mathfrak{f}^{\nabla}_{\circ},a_{\circ}/p^{N})$ (4.14)
@@
[…]
@@
First, we shall prove assertion (i). Denote by $\ell_{\circ}:=(\ell_{\circ}^{[1]},\cdots,\ell_{\circ}^{[m_{\circ}]})$ the type of $\mathfrak{f}_{\circ}$. The assertion is of local nature, so $\mathscr{F}:=(\mathcal{F},\nabla)$ may be identified with a direct sum of $\mathscr{O}_{d,\flat}$’s (cf. the discussion around (3.19)). Since $(\mathcal{F},\nabla)$ has parabolic weights $a_{0}$, the equality (3.18) implies that the problem is reduced to the situation where $\mathscr{F}$ satisfies the following conditions:
@@
  • For each $j=0,\cdots,m_{\circ}$, the equality $\mathcal{F}_{\circ}^{[j]}=\bigoplus_{j^{\prime}=1}^{j}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}$ holds. Also, the projection $\mathfrak{f}_{\circ}^{[j]}$ coincides with the projection $\mathcal{F}_{\circ}^{[j]}$ given by forgetting the last direct summand.
@@
($j=0,\cdots,m_{\circ}$), where $D^{(N)}_{\circ}:=\mathrm{Im}(\sigma_{\circ}^{(N)})$. Hence, the equality $\mathcal{F}^{\nabla[j]}_{\circ}=\bigoplus_{j^{\prime}=1}^{j}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}$ holds and $\mathfrak{f}_{\circ}^{\nabla[j]}:\mathcal{F}^{\nabla[j]}_{\circ}\twoheadrightarrow\mathcal{F}^{\nabla[j-1]}_{\circ}$ coincides the projection given by forgetting the last direct summand. Thus, $\mathscr{F}^{\nabla}$ specifies an object in $\mathcal{B}un_{\mathscr{U}^{(N)},a_{0}/p^{N}}$, and this completes the proof of assertion (i).
@@
Since the formation of $\mathscr{F}_{\flat}\mapsto\mathscr{F}^{\nabla}$ is verified to be functorial, this operation yields a functor $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}\rightarrow\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}$.